OpenStudy (anonymous):

A2+B2=C2 A=5 B=2 C=? Can someone solve for me?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (mathteacher1729):

Is this saying "a times 2 plus b times 2 plus c times 2?"

OpenStudy (anonymous):

c=7

OpenStudy (anonymous):

2 is saying that its squared

OpenStudy (mathteacher1729):

Ah, ok. \(a^2+b^2=c^2\) where \(a=5,b=2\) and we must solve for \(c\). So this means \[(5)^2+(2)^2 = c^2\] What is 5 squared plus 2 squared?

OpenStudy (anonymous):

c = \[\sqrt{29}\]

OpenStudy (anonymous):

um squared is like multipling it by itself rig ht

OpenStudy (mathteacher1729):

Yes. Squared = multiply the number by itself. :) So 4 squared = 4*4 = 16 And 6 squared = 6*6 = 36 etc.

OpenStudy (anonymous):

so 5 squared is 25 and 2 squared is 4